OpenStudy (anonymous):

2/3,3/4,4/5,5/6,6/7,.... which does it use n/n+1 or n-1/n, or n+1/n+2????

OpenStudy (anonymous):

makes no difference at all

OpenStudy (anonymous):

i have to choose one as the answer

OpenStudy (anonymous):

if you write \(\frac{n}{n+1}\) then \(n\) starts at 2 if you write \(\frac{n+1}{n+2}\) then \(n\) starts at 1 if you write \(\frac{n+2}{n+3}\) then \(n\) starts at 0

OpenStudy (anonymous):

can you write the exact question? because as stated it makes no difference which you use, depends only on where you start

OpenStudy (anonymous):

Which of the following represents the general term for the sequence given?

OpenStudy (anonymous):

thats what it wants to know for the problem

jimthompson5910 (jim_thompson5910):

I would go with the formula that starts at n = 1 since that's usually what you start with

OpenStudy (anonymous):

i would tell your teacher that the problem is not written by a mathematician

OpenStudy (anonymous):

(ok, probably not a good idea)

OpenStudy (anonymous):

its online school